OpenStudy (abstracthaven):

write the equation in slope-intercept form and standard form. through(-2,4) and (1,3)

OpenStudy (anonymous):

To write the equation in standard form, you must first find the slope. By the slope formula, that is (3-4)/(1+2)=-1/3. next, you recognize that the equation for a line in slope-intercept form is y-y1=m*(x-x1). Thus, all you need to do is plug in the values. Namely, y-3=-1/3*(x-1). For slope intercept form, you just take what you learned from standard form but plug in one more thing. y=-1/3*x+b. To solve for b, set y equal to 3 and x equal to 1. This causes b to be 4/3. thus, the equation in slope-intercept form is y=-1/3*x+4/3

OpenStudy (anonymous):

You have to first find your slope using the formula \[m=\frac{x_2-x_1}{y_2-y_1}\] And then use one of the points listed to substitute m, and a respect x and y to find your equations... y-y1=m(x-x1) y=mx+c is slope-intercept form and ax+by-c=0 is standard form but some books describe standard as the same as general form or differently... and general form is ax+by=c so refer to your textbook or lecturer notes to see how they describes it.

OpenStudy (abstracthaven):

thank you i needed formulas :D

OpenStudy (anonymous):

You're welcome!